40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1509     Entry time: Thu Apr 23 16:27:24 2009
Author: Yoichi 
Type: Update 
Category: Locking 
Subject: Locking with the cryo-pump 
The last night, the IFO was unstabler than usual and the locking script often failed before reaching the power up stage.
The failure happened at random points.
I'm not sure if this is related to the operation of the cryo-pump.
The mode cleaner reflection image seemed to move around more than usual. Maybe it was just a high seismic night.
ELOG V3.1.3-